OpenStudy (anonymous):

a student claims that the best way to solve the system at the right is by substitution.Do you agree? explain system:y-3x=4 y-6x=12

OpenStudy (anonymous):

in stuck on this question

OpenStudy (mathmale):

Either of the given equations is very easily solved for y. What does that tell you?

OpenStudy (anonymous):

umm....that it can be substitution?

OpenStudy (texaschic101):

yes....you can easily solve one of the equations for y, then sub that back into the other equation. Do you need me to show you ?

OpenStudy (anonymous):

YES! plz it would be a big help. @texaschic101

OpenStudy (texaschic101):

y - 3x = 4......solve for y by adding 3x to both sides y = 3x + 4 now sub 3x + 4 in for y in the other equation to find x y - 6x = 12 (3x + 4) - 6x = 12 3x - 6x = 12 - 4 -3x = 8 x = -8/3 now you would sub -8/3 in for y in either of the original equations y - 3x = 4 y - 3(-8/3) = 4 y + 24/3 = 4 y = -24/3 + 4 y = -24/3 + 12/3 y = -12/3 y = -4